# An Exact Solution Algorithm for the Bi-Level Optimization Problem of Electric Vehicles Charging Station Placement

### v1

@@ 3   Extract duals $\sigma_{rs},\zeta_{ij},\gamma_{l}$ over $\bar{\Pi}$
5      Build a BCSPP with arc‐costs $\zeta_{ij}$ and station‐costs $\gamma_{l}$
6      Solve BCSPP to get path $\pi^{\star}_{rs}$ minimizing $\sum\zeta_{ij}-\sum\gamma_{l}$.
@@ 7      Compute its reduced cost:
$c^{\star}\;=\;-\sigma_{rs}\;+\;\sum_{(i,j)\in\pi^{\star}_{rs}}\zeta_{ij}\;-\;\sum_{l\in\pi^{\star}_{rs}}\gamma_{l}$
$\text{RC}_{\min}\leftarrow\min(\text{RC}_{\min},\,c^{\star})$.
8    if $c_{\pi^{\star}_{rs}}<0$ then
@@ 6 Numerical Results
We conducted numerical experiments on three different in size networks. The Eastern Massachusetts network, which comprises 74 zones, 74 nodes, and 248 links, and the Anaheim network consists of 38 zones, 416 nodes, and 914 links. In addition, various number of new charging station candidate locations were introduced. To accurately model EV routing behavior, we expanded the original network of Eastern Massachusetts by using battery state considerations. The expansion network resulted in an expanded network of 1,624 nodes and 2,612 links. The Anaheim network resulted in 4,015 nodes and 6,191 links in expanded form. The next larger network we have tried our algorithm on is the Barcelona transportation network. This network has 110 zones, 1,020 nodes, and 2,512 links. The network was expanded to 7,559 nodes and 9,169 links. This expanded representation allowed route choices that respect battery constraints and charging requirements. All computational experiments were performed on a MacBook Air equipped with an Apple M3 chip and 8 GB of unified memory, running macOS.
@@
We derive the $t_{ij}(x_{ij})$, which is the link flow in our model, from the BPR function, which takes the form:
$t_{ij}(x_{ij})=t_{ij}^{0}\cdot\left(1+\alpha~~\left(\frac{x_{ij}}{c_{ij}}\right)^{\beta}\right)$ (22)
@@
$\alpha$ and $\beta$ are shape parameters that can be calibrated to data. It is common to use $\alpha=0.15$ and $\beta=4$. $t_{ij}^{0}$ is the free flow travel time on link $(i,j)$, and $c_{ij}$ is the capacity of link $(i,j)$. This equation is used to derive the travel time of a link $(i,j)$. Figures 2 and 3 illustrate the convergence behavior and optimality gap percentage of our BPC algorithm on the Eastern Massachusetts network, respectively. These graphs are for the case of 30 candidate locations for charging stations. The convergence graph demonstrates the progression of the LB and UB. The gap between the two bounds decreases steadily until convergence is achieved. The last B&B node gap is 0.79%. Specifically, the optimal solution was identified after evaluating 26 nodes in the B&B tree, which shows the effectiveness of using value function cuts. Figure 4 shows the number of paths in each B&B node, indicating the success of CG in identifying and adding useful paths.
@@
Table 1 presents a detailed breakdown of computational performance across different problem sizes on the Eastern Massachusetts network, where ”Cand.” shows the number of candidate locations. The ”MTAP (s)” column shows the time required to solve the MTAP, which determines traffic flow patterns given the current network configuration. The ”CG (s)” represents the time needed to identify new paths to add to the HPR. ”HPR (s)” shows the time spent solving HPR using CPLEX Solver [cplex2009v12]. ”Total (s)” indicates the total algorithm termination time, ”VF cuts” represents the total number of valid function cuts generated throughout the solution process, ”Nodes” indicates the B&B number of nodes until termination, and ”Comb.” displays the total number of feasible combinations under the budget constraint. ”Gap %” shows the gap in the last node, and ”Obj” shows the optimal objective function value.
Table 1: Computational Performance Results on Eastern Massachusetts
Cand. Nodes Comb. Gap (%) Obj. Total (s) MTAP (s) CG (s) HPR (s) VF cuts
5 3 24 0.0 18633.726 25.185 4.653 1.610 0.447 2452
10 8 347 0.166 18633.726 35.003 2.513 9.740 1.571 2481
20 8 10480 0.767 17052.465 69.720 12.144 26.917 4.823 2534
25 10 54420 0.0 16126.790 77.773 11.733 32.894 5.123 2528
30 26 304786 0.798 15896.640 205.526 55.378 95.578 19.813 2715
40 24 $\infty$ 0.0 8271.450 1797.142 986.601 601.784 164.271 2836
The results reveal that the MTAP component constitutes the primary computational bottleneck, consuming the largest portion of total runtime and showing the most considerable growth with problem size. This indicates that employing faster TAP algorithms could highly improve overall performance. The CG shows moderate scaling, while the HPR solution by CPLEX remains highly efficient even for complex instances. The algorithm successfully handles problems with over 300,000 feasible combinations, maintaining reasonable iteration counts. Table 2 shows the computation performance on the Anaheim network.
Table 2: Computational Performance Results on Anaheim
Cand. Nodes Comb. Gap (%) Obj. Total (s) MTAP (s) CG (s) HPR (s) VF cuts
20 10 10480 0.001 40034.745 190.027 9.628 70.576 6.116 5629
25 20 54420 0.0 40030.506 479.680 13.911 253.716 23.206 6924
30 36 304786 0.299 40030.847 957.177 49.112 617.631 85.924 7208
40 64 $\infty$ 1.622 39840.151 3719.143 159.491 2553.716 872.949 7336
We will stop the solution process if the total calculated time after a step is more than 1 hour. The last row of Table 2 shows the total termination time of 3719.14 s, which we stopped on and report the existing gap. The gap is 1.622%, which shows the superiority of our algorithm that can achieve this gap on a large and realistic network in a very fast runtime. The time for MTAP is relatively small since we have not we have not explore all the branches yet, so the times MTAP was running after finding a fixed solution is lower compared to previous rows. The most time consuming part of the algorithm for this larger network, compared to Eastern Massachusetts, is the CG. This means that the number of paths added to the network is high.
Table 3 shows the result on the Barcelona network. The run time is higher compared to the Anaheim network since the network is larger. Similar to Anaheim, we put a one hour threshold. When we choose 40 candidates for the charging station, after 3,668.18 (s), the solution process stopped with a 3.8 % gap.
Table 3: Computational Performance Results on Barcelona
Cand. Nodes Comb. Gap (%) Obj. Total (s) MTAP (s) CG (s) HPR (s) VF cuts
20 10 10480 0.0 36130.354 742.515 25.330 111.271 11.710 9442
25 16 54420 0.032 36010.367 909.689 31.016 199.218 17.597 9638
30 14 304786 0.961 36109.161 1080.887 79.951 252.499 17.725 9715
40 35 $\infty$ 3.802 35160.987 3668.182 94.790 2094.545 150.110 9738
@@
Our proposed algorithm demonstrates exceptional computational efficiency compared to existing approaches in the literature. In a study by 10034431, their proposed heuristic algorithm required 2.3 hour for reaching 4.5 % gap and their benchmark exact solution algorithm required 173.1 hours to converge on a hypothetical network of 12 nodes and 32 arcs. In another study, 9690622 reached the convergence of upper and lower bounds with a gap of 4.58%, and the CPU time for their algorithm was 21.4 hours on a hypothetical network dataset including 18 nodes and 58 links. In contrast, our algorithm consistently delivers exact optimal solutions within minutes rather than hours with guaranteed optimality. In addition, in another study of charging station location by doi:10.1287/trsc.2021.0494, they stated that for the large network of more than 300 nodes, Gurobi was unable to obtain even an incumbent solution within a two-hour time limit. So our result represents a computational speedup of over two orders of magnitude compared to existing methods, while simultaneously providing superior solution quality through exact optimization rather than heuristic approximations. The efficiency comes from our cutting plane generation strategy, and our effective HPR solution via CPLEX and the CG algorithm, which starts just with the useful paths and then adds as needed.

### v2

@@ 3   Extract duals $\sigma_{rs},\zeta_{ij},\gamma_{l}$ over $\bar{\Pi}$
5      Build a BCSPP with arc‐costs $\zeta_{ij}$, station‐flow‐costs $\gamma_{l}$, and charging‐costs $c\cdot\gamma_{l}^{\pi}$
6      Solve BCSPP to get path $\pi^{\star}_{rs}$ minimizing $\sum\zeta_{ij}-\sum\gamma_{l}+\sum c\cdot\gamma_{l}^{\pi}$.
@@ 7      Compute its reduced cost:
$\bar{c}^{\star}\;=\;-\sigma_{rs}\;+\;\sum_{(i,j)\in\pi^{\star}_{rs}}\zeta_{ij}\;-\;\sum_{l\in\pi^{\star}_{rs}}\gamma_{l}\;+\;c^{\pi^{\star}_{rs}}$
$\text{RC}_{\min}\leftarrow\min(\text{RC}_{\min},\,\bar{c}^{\star})$.
8    if $\bar{c}_{\pi^{\star}_{rs}}<0$ then
@@ 6 Numerical Results
We conducted numerical experiments on three different networks. The Eastern Massachusetts network, which comprises 74 zones, 74 nodes, and 248 links, and the Anaheim network consists of 38 zones, 416 nodes, and 914 links. In addition, various number of new charging station candidate locations were introduced. To accurately model EV routing behavior, we expanded the original network of Eastern Massachusetts by using battery state considerations. The expansion network resulted in an expanded network of 1,624 nodes and 2,612 links. The Anaheim network resulted in 4,015 nodes and 6,191 links in expanded form. The next larger network we have tried our algorithm on is the Barcelona transportation network. This network has 110 zones, 1,020 nodes, and 2,512 links. The network was expanded to 7,559 nodes and 9,169 links. This expanded representation allowed route choices that respect battery constraints and charging requirements. All computational experiments were performed on a MacBook Pro equipped with an Apple M4 Pro chip and 64 GB of unified memory, running macOS.
@@
We derive the $t_{ij}(x_{ij})$, which is the link travel time in our model, from the BPR function, which takes the form:
$t_{ij}(x_{ij})=t_{ij}^{0}\cdot\left(1+\alpha\left(\frac{x_{ij}}{C_{ij}}\right)^{\beta}\right)$ (29)
$\alpha$ and $\beta$ are shape parameters that can be calibrated to data. It is common to use $\alpha=0.15$ and $\beta=4$. $t_{ij}^{0}$ is the free flow travel time on link $(i,j)$, and $C_{ij}$ is the capacity of link $(i,j)$. This equation is used to derive the travel time of a link $(i,j)$. Similarly, the charging station congestion function $t_{l}(v_{l})$ follows a BPR-style formulation:
$t_{l}(v_{l})=t_{l}^{0}\cdot\left(\alpha\left(\frac{v_{l}}{K_{l}}\right)^{\beta}\right)$ (30)
where $K_{l}$ is the service capacity of station $l$ representing the maximum flow rate the station can handle efficiently. We use the same shape parameters $\alpha=0.15$ and $\beta=4$ for consistency with link travel time. The station capacity is determined by the investment cost through the relation $K_{l}=\kappa\cdot C_{l}$, where $\kappa$ is a capacity cost proportionality constant and $C_{l}$ is the installation cost at station $l$. Note that this equation captures the delay at the station due to congestion and queue, not the charging time. Charging time was captured on the third component of the objective function, which varied based on the arrival battery state. This formulation reflects the realistic assumption that stations with higher investment have greater capacity and can serve more vehicles with less congestion. The charging price is set to $c$ per unit time (e.g., dollars per hour), so the path-dependent charging cost is $c^{\pi}=\sum_{l\in\mathcal{C}}c\cdot\gamma_{l}^{\pi}$, where $\gamma_{l}^{\pi}$ is the charging time for a vehicle on path $\pi$ at station $l$ based on its battery state upon arrival. This monetary cost captures both the direct electricity cost and the opportunity cost of time spent charging, ensuring that users with different arrival battery states incur different costs proportional to their required charging duration.
The charging time $\gamma_{l}^{\pi}$ is not an exogenous input but is endogenously determined through our battery-expanded network representation. In the expanded network, each node is represented as a tuple $(i,b)$, where $i$ is the physical location and $b$ is the remaining battery level. As vehicles traverse the network, the battery state is tracked at each node along the path. When a vehicle following path $\pi$ arrives at charging station $l$, the expanded network directly provides the arrival battery level $b_{l}^{\pi}$. The charging time is then calculated as:
$\gamma_{l}^{\pi}=(B^{\max}-b_{l}^{\pi})\cdot\rho$ (31)
where $B^{\max}$ is the full battery capacity, $b_{l}^{\pi}$ is the arrival battery level at station $l$ for a vehicle following path $\pi$, and $\rho$ is the charging rate (time required to charge one unit of battery, e.g., hours per percentage point). Consequently, the charging cost for a vehicle on path $\pi$ at station $l$ becomes:
$c^{\pi}\cdot\gamma_{l}^{\pi}=(c+\tau)\cdot(B^{\max}-b_{l}^{\pi})\cdot\rho$ (32)
This formulation ensures that vehicles arriving with depleted batteries spend more time charging and more cost and thus incur higher perceived charging costs than those arriving with higher charge levels, capturing the path dependent nature of charging costs in our model. The total charging cost $c^{\pi}$ that depends on path aggregates these costs across all charging stations visited along path $\pi$.
In the following numerical experiments, we assume that EVs depart from their origins with a full battery, i.e., $b=B^{\max}$. To ensure that the charging station placement problem remains practically relevant, we scale the link lengths in the network to represent long distance travel scenarios where en-route charging becomes necessary. This scaling reflects the intended application of our model for strategic infrastructure planning on highway networks or inter city corridors, where trip distances typically exceed the driving range of EVs. It is important to note that the full battery assumption is not a fundamental limitation of our methodology. The battery-expanded network representation is flexible and can accommodate any initial battery state. Specifically, the model can be extended to handle: (i) a fixed partial initial battery for all users (e.g., $b=0.7\cdot B^{\max}$), representing conservative planning scenarios; (ii) origin destination specific initial battery states $b_{rs}^{\text{init}}$, which can show the heterogeneous trip characteristics; or (iii) stochastic initial battery states drawn from a distribution to reflect real world variability in departure conditions. These extensions require only modifying the initialization of the battery state in the expanded network construction, while the core algorithmic framework, including the BPC algorithm, value function cuts, and column generation remains entirely unchanged. This flexibility ensures that our approach is applicable to a wide range of practical scenarios beyond the specific assumptions used in our numerical experiments.
The model parameters were set as follows. The base station cost was set to $C=200$ monetary units, with individual station costs randomly drawn from the range $[0.1C,1.0C]$ using a uniform distribution which reflects heterogeneity in installation costs across candidate locations. The budget was calculated as a proportion of the total cost of all candidate stations, i.e., $B=B_{\text{prop}}\times\sum_{l\in\mathcal{L}}c_{l}$, where $B_{\text{prop}}$ represents the budget proportion varied across experiments. Station capacity was determined proportionally to investment cost using the relationship $K_{l}=\kappa\cdot c_{l}$, where the capacity-cost ratio $\kappa=0.1$ ensures that stations with higher installation costs have correspondingly higher service rates. The charging price was set to $p=10$ per unit of flow, representing the revenue collected from users charging their vehicles. The value of time was set to $\tau=1.0$ \$/minute (equivalent to \$60/hour), a commonly used value in transportation economics. The charging time per unit of battery was set to $\gamma=1.0$ minute, meaning that recharging from empty to full (100 units) requires 100 minutes at the base rate. These parameter values were chosen to reflect realistic operating conditions while maintaining computational tractability.
@@
Figures 2 illustrate the convergence behavior and optimality gap percentage of our BPC algorithm on the Anaheim network, respectively. These graphs are for the case of 30 candidate locations for charging stations. The convergence graph demonstrates the progression of the LB and UB. The gap between the two bounds decreases steadily until convergence is achieved. Specifically, the optimal solution was identified after evaluating 33 nodes in the B&B tree, achieving a 0.00% optimality gap, which shows the effectiveness of using value function cuts. Figure 3 shows the number of paths in each B&B node, indicating the success of CG in identifying and adding useful paths.
@@
Table 2 presents a detailed breakdown of computational performance across different problem sizes on all three networks, where “Candidates” shows the number of candidate locations. The “TAP (s)” column shows the time required to solve the MTAP, which determines traffic flow patterns given the current network configuration. The “Pricing (s)” represents the time needed to identify new paths to add to the HPR. “RMP (s)” shows the time spent solving HPR using CPLEX Solver [3]. “Total (s)” indicates the total algorithm termination time, “VF cuts” represents the total number of value function cuts generated throughout the solution process, “BB_Nodes” indicates the B&B number of nodes until termination, and “Comb.” displays the total number of feasible combinations under the budget constraint. “Gap (%)” shows the optimality gap at termination, “Objective_UB” shows the upper bound objective function value, “Revenue” shows the charging revenue collected, and “Unmet_Demand” shows the penalty for unserved demand.
Table 2: Summary of computational results across all networks
Candidates Comb. BB_Nodes Gap (%) Objective_UB Revenue Unmet_Demand Num_Selected Num_Paths Total time(s) TAP (s) RMP (s) Pricing(s) VF_cuts
EasternMassachusetts ($B_{\text{Prop}}=0.4$)
5 24 7 0.00 17142.60 59.16 172.62 2/5 568 18.76 2.84 1.49 0.54 2409
10 347 9 0.00 15753.97 239.00 162.25 4/10 2012 61.30 41.43 4.04 1.19 2478
20 10480 2 0.60 15300.04 278.20 155.78 7/20 3835 80.00 59.92 4.12 0.63 2617
25 54420 29 0.00 13954.52 428.28 145.77 10/25 6068 101.56 60.04 18.99 6.80 2630
30 304786 47 0.00 13411.06 417.12 140.45 13/30 10904 235.99 150.16 48.08 20.31 2848
Anaheim ($B_{\text{Prop}}=0.3$)
20 10480 19 0.47 40225.16 79.26 409.24 6/20 1968 281.54 47.34 174.98 5.23 5599
25 54420 17 0.75 40031.32 23.30 411.95 7/25 3672 427.32 71.24 258.16 8.61 6983
30 304786 33 0.00 40225.16 95.67 488.66 8/30 4091 779.44 124.97 535.82 16.70 7076
40 ¿ $10^{6}$ 93 0.00 39885.12 181.12 406.82 12/40 16174 3169.09 276.80 2521.39 227.03 7908
Barcelona ($B_{\text{Prop}}=0.3$)
10 10480 9 0.60 36245.05 27.88 360.56 3/10 141 856.27 189.16 383.52 4.84 9303
20 54420 28 0.15 35990.40 52.95 359.90 7/20 199 1674.98 236.21 1106.95 12.18 9435
30 304786 25 0.77 35985.51 173.52 358.82 9/30 785 2325.35 294.87 1641.82 17.06 9723
40 ¿ $^{6} 101 0$.91 35506.57 299.43 354.82 13/40 6000 10530.83 395.28 9597.46 112.89 9928
For the Eastern Massachusetts network, the algorithm achieves 0.00% optimality gaps for most instances, with only the 20-candidate instance terminating with a small gap of 0.60%. The largest instance (30 candidates, over 300,000 feasible combinations) is solved to optimality in 235.99 seconds. The results reveal that the TAP component constitutes the primary computational bottleneck, growing from 2.84 seconds for 5 candidates to 150.16 seconds for 30 candidates. The RMP and Pricing (CG) components also show considerable growth with problem size. The algorithm successfully identifies optimal solutions with up to 13 out of 30 stations selected.
For the Anaheim network, the algorithm solves all instances, with the two smaller instances (20 and 25 candidates) terminating with gaps of 0.47% and 0.75%, respectively, while the larger instances (30 and 40 candidates) achieve 0.00% gaps. The largest Anaheim instance with 40 candidates required 3169.09 seconds and explored 93 B&B nodes, selecting 12 out of 40 candidate stations. The most time-consuming component for this network is the RMP, consuming 2521.39 seconds out of the total runtime for the 40-candidate instance. This means that the number of paths added to the network is high, increasing the complexity of the master problem.
For the Barcelona network, which is the largest network tested, the algorithm maintains optimality gaps below 1% across all instances. The most challenging instance with 40 candidates required 10530.83 seconds (approximately 2.9 hours) and explored 101 B&B nodes, achieving a 0.91% gap while selecting 13 out of 40 stations. The RMP dominates the computation time, accounting for 9597.46 seconds of the total runtime. This indicates that employing faster TAP algorithms and more efficient master problem solvers could considerably improve overall performance.
@@
Our proposed algorithm demonstrates exceptional computational efficiency compared to existing approaches in the literature. In a study by Mirheli and Hajibabai [20], their proposed heuristic algorithm required 2.3 hours for reaching 4.5% gap, and their benchmark exact solution algorithm required 173.1 hours to converge on a hypothetical network of 12 nodes and 32 arcs. In another study, Mirheli and Hajibabai [19] reached the convergence of upper and lower bounds with a gap of 4.58%, and the CPU time for their algorithm was 21.4 hours on a hypothetical network dataset including 18 nodes and 58 links. In contrast, our algorithm consistently delivers optimal or near-optimal solutions within minutes to hours with guaranteed optimality bounds. In addition, in another study of charging station location by Kınay et al. [13], they stated that for the large network of more than 300 nodes, Gurobi was unable to obtain even an incumbent solution within a two-hour time limit. Based on published results from the literature, our algorithm solves substantially larger networks with tighter optimality gaps in comparably less time than existing methods. While a direct head-to-head comparison on identical instances is not available due to differences in problem formulations and the unavailability of prior implementations, the magnitude of the difference is notable: existing exact methods require tens to hundreds of hours on networks with fewer than 50 nodes, whereas our algorithm handles networks with over 1,000 nodes within hours. The efficiency comes from our cutting plane generation strategy, and our effective HPR solution via CPLEX and the CG algorithm, which starts just with the useful paths and then adds as needed.
7 Sensitivity Analysis
In this bi-level optimization model for EV charging station placement, the budget parameter is the only factor that meaningfully influences the optimal station configuration. This occurs because the model exhibits captive demand in which EV users whose battery is insufficient to complete their trip must charge regardless of waiting time, congestion delays, or charging price. Consequently, both components of the objective function (maximizing revenue and minimizing unmet demand) are aligned rather than competing: building more stations always improves both metrics simultaneously. Changing the weight parameter ($w$) only rescales the penalty for unmet demand but does not alter which stations are selected, since all feasible configurations rank in the same order regardless of $w$. Similarly, varying the kappa parameter (capacity scaling factor) affects congestion delays but not station selection, because users have no alternative, and they will use the nearest available station even with high waiting times. The charging price also has no effect on station configuration because demand is inelastic; users cannot choose to forgo charging when their battery is depleted. Therefore, the budget constraint serves as the true limiting factor, and sensitivity analysis is most meaningful when examining how different budget levels affect which stations are selected and the resulting network performance.
7.1 Budget Sensitivity Analysis
To investigate the impact of budget availability on optimal charging station deployment, we conducted a sensitivity analysis on the Eastern Massachusetts network with 30 candidate locations, varying the budget proportion $B_{\text{prop}}$ from 0.1 to 0.9.
Table 3: Sensitivity analysis on budget proportion for Eastern Massachusetts (30 candidates)
$B_{\text{prop}}$ Gap (%) Num_Selected BB_Nodes Num_Paths Total(s) TAP(s) RMP(s) Pricing(s) Budget_Util (%)
0.1 0.00 4/30 11 4495 86.64 23.11 47.15 3.31 99.80
0.2 0.22 6/30 15 5473 116.48 26.29 71.90 4.76 97.61
0.3 0.00 9/30 27 4415 167.67 26.35 119.17 8.48 96.40
0.4 0.00 12/30 1 3794 52.12 27.57 10.29 0.50 91.45
0.5 0.00 12/30 1 3794 51.09 26.89 10.23 0.50 77.78
0.6 0.00 12/30 1 3794 51.46 27.08 10.15 0.48 65.60
0.7 0.00 12/30 1 3794 51.08 26.93 10.06 0.47 65.33
0.8 0.00 12/30 1 3794 51.40 27.17 10.10 0.47 49.03
0.9 0.00 12/30 1 3794 51.42 27.36 9.89 0.48 39.13
Figure 4: Branch-and-bound nodes and computation time versus budget proportion
Figure 5: Number of selected stations versus budget proportion
As shown in Figure 5, the number of selected stations increases monotonically from 4 at $B_{\text{prop}}=0.1$ to 12 at $B_{\text{prop}}=0.4$, after which it plateaus. This saturation indicates that 12 stations are sufficient to serve all feasible demand in the network, and additional budget beyond $B_{\text{prop}}=0.4$ provides no further benefit in terms of station deployment. This finding is practically significant for planners, as it identifies the minimum budget threshold beyond which additional investment yields no improvement in network coverage.
The computational behavior exhibits a non-monotonic pattern with respect to budget, as illustrated in Figure 4. The number of B&B nodes peaks at 27 for $B_{\text{prop}}=0.3$ and drops sharply to just 1 node for $B_{\text{prop}}\geq 0.4$. Correspondingly, the total computation time reaches its maximum of 167.67 seconds at $B_{\text{prop}}=0.3$ and decreases to approximately 51 seconds for higher budget levels. This behavior can be explained by the combinatorial difficulty of the problem: when the budget is tight, the algorithm must explore more branches to determine the optimal subset of stations under binding budget constraints. At moderate budgets ($B_{\text{prop}}=0.3$), the problem reaches peak difficulty because there are many near-optimal feasible combinations to evaluate. Once the budget is sufficiently large ($B_{\text{prop}}\geq 0.4$), the budget constraint becomes non-binding, and the algorithm can identify the optimal solution at the root node without branching.
The budget utilization decreases steadily from 99.80% at $B_{\text{prop}}=0.1$ to 39.13% at $B_{\text{prop}}=0.9$, confirming that higher budget levels lead to increasingly slack budget constraints. These results demonstrate that our BPC algorithm adapts efficiently to varying problem difficulty, requiring more computational effort when the problem is combinatorially challenging (tight budgets) and converging rapidly when the solution structure is simpler (loose budgets).
7.2 Station Capacity Sensitivity Analysis
We examine the effect of the capacity-cost proportionality constant $\kappa$ on charging station congestion and optimal deployment. Recall that the service capacity of each station is determined by $K_{l}=\kappa\cdot C_{l}$, where $C_{l}$ is the installation cost. Higher values of $\kappa$ yield greater station capacity, reducing congestion delays experienced by EV users. We vary $\kappa\in\{0.2,0.5,1.0,1.5\}$ on the Eastern Massachusetts network with 20 candidate locations and $B_{\text{prop}}=0.3$.
Table 4 summarizes the station capacities for representative stations at each $\kappa$ value, illustrating how low $\kappa$ values result in station capacities well below the expected flow levels.
Table 4: Station capacity $K_{l}$ for selected stations at different $\kappa$ values
Station Cost ($C_{l}$) $K$ ($\kappa$=0.2) $K$ ($\kappa$=0.5) $K$ ($\kappa$=1.0) $K$ ($\kappa$=1.5)
231 196.90 3.96 9.90 19.80 29.70
226 183.75 3.70 9.24 18.48 27.72
182 112.03 2.25 5.63 11.27 16.90
187 92.89 1.87 4.67 9.34 14.01
179 66.61 1.34 3.35 6.70 10.05
Consequently, small changes in $\kappa$ can produce large differences in congestion when stations operate near or above capacity.
Table 5: Sensitivity of station congestion to $\kappa$ on Eastern Massachusetts (20 candidates, $B_{\text{prop}}=0.3$)
$\kappa$ Num_Selected Avg Delay (min) Revenue Gap (%) Total(s)
0.2 6/20 24.66 284.57 0.00 293.57
0.5 6/20 20.25 284.57 0.00 307.58
1.0 6/20 1.36 284.57 0.00 294.44
1.5 5/20 0.66 284.57 0.90 172.47
The results reveal that $\kappa$ has a pronounced effect on station congestion while leaving the overall network design largely unchanged. For $\kappa=0.2$, the average congestion delay across active stations is 24.66 minutes, reflecting severely constrained capacity where station flows substantially exceed the effective service rate. As $\kappa$ increases to 0.5, the delay decreases modestly to 20.25 minutes. However, at $\kappa=1.0$, the delay drops dramatically to 1.36 minutes, representing a reduction of over 94% compared to $\kappa=0.2$.
Notably, the same set of 6 stations is selected for $\kappa\in\{0.2,0.5,1.0\}$, with identical revenue of 284.57, indicating that the optimal station locations are robust to capacity assumptions within this range. At $\kappa=1.5$, one fewer station is selected (5 instead of 6), suggesting that when individual stations have sufficiently high capacity, fewer facilities are needed to serve the same demand. The revenue remains unchanged at 284.57, as the same total flow is served regardless of the number of stations. The computation time also decreases at $\kappa=1.5$ (172.47 seconds versus approximately 300 seconds for other values), consistent with the reduced combinatorial complexity of selecting fewer stations.
These findings have practical implications for infrastructure planning. The parameter $\kappa$ reflects the relationship between investment cost and service capacity, which depends on technology choices such as the number of chargers per station and their power output. Our results suggest that investing in higher-capacity stations (larger $\kappa$) can substantially reduce user waiting times without altering the optimal spatial distribution of stations, up to a threshold beyond which fewer but larger stations become preferable.